### v2

Abstract
Over the past six years, molecular transformer models have become part of the computational toolbox for drug design tasks. Most of the current transformer-based chemical language models are pre-trained on millions to billions of molecules from unlabeled datasets like ZINC, or ChEMBL. However, the improvement from such scaling in dataset size is not confidently linked to improved molecular property prediction.
The aim of this study was to investigate and overcome some of the limitations of transformer models in predicting molecular properties. Specifically, we examine the impact of pre-training dataset size and diversity on the performance of transformer models, and investigate the use of domain adaptation as a technique for improving model performance. First, our findings indicate that increasing pre-training dataset size beyond $\sim 400K$ molecules (i.e., 30%) from the pre-training dataset does not result in a significant improvement on four ADME endpoints, namely, solubility, permeability, microsomal stability, and plasma protein binding. Further analysis indicates that the limited benefits of increasing the pre-training dataset size may be attributed to its lower diversity in Murcko scaffolds compared to the downstream datasets. Second, our results demonstrate that using domain adaptation by further training the transformer model on a small set of domain-relevant molecules, i.e., a few hundred to a few thousand, using multi-task regression of physicochemical properties was sufficient to significantly improve performance for three out of the four investigated ADME endpoints (P-value $<0.001$). Finally, we observe that a model pre-trained on $\sim 400$K molecules and domain adopted on a few hundred/thousand molecules perform similarly (P-value $>0.05$) to more complicated transformer models like MolBERT [1] (pre-trained on $\sim 1.3$M molecules) and MolFormer [2] (pre-trained on $\sim 100$M molecules). A comparison to a random forest model trained on basic physicochemical properties showed similar performance to the examined transformer models.
We believe that current transformer models can be improved through further systematic analysis on pre-training and downstream data, pre-training objectives, and scaling laws, ultimately leading to better and more helpful models.
footnotetext: $\dagger$ Co-first authors.footnotetext: §$\S$ Co-corresponding authors.
1 Introduction
Molecular property prediction (MPP) is a task at the heart of diverse cheminformatics and drug design challenges. Molecular properties can range from inherent physical features of the molecule like lipophilicity or solubility to more complex results of the physiological properties like toxic effects of the molecule on an organism [3]. Supervised learning (SL) has been used to map predefined or heuristic molecular descriptors to such properties [3, 4]. However, freely available datasets for MPP tasks usually consist of only a few hundred to a few thousand molecules [3, 5] due to the complex and expensive experimental processes to generate the data [6]. Existing property prediction methods suffer from limitations regarding data representation. On the one end, human-made descriptors like predefined fingerprints require expert knowledge and are restricted to known rules or patterns [7]. On the other end, data-driven methods like deep learning require a large amount of labeled data [8].
Self-supervised learning (SSL) has been used as an alternative to learning from labeled data as in supervised learning [9]. In SSL, the model is initially trained on large unlabeled data to learn intrinsic relationships within the input. These relationships can be obtained by using tasks like recovering parts of the input using information from other parts of the input [10, 11]. Such an SSL model is then thought of as a foundation model that can be generalized to different downstream tasks [11, 12]. The past decade has seen a breakthrough in the field of Natural Language Processing (NLP) with the introduction of the SSL-based transformer model [13], which has inspired multiple works to adopt similar schemes for sequence-based representations of molecules [14, 15, 16].
The transformer model [13] is a sequence-to-sequence model composed of an encoder-decoder architecture and trained on the next token prediction (NTP) objective. In this objective, the model is optimized such that the decoder model correctly predicts the next token (i.e., subword) given the previous tokens. While the transformer model was built for machine translation, BERT (Bidirectional Encoder Representations from Transformers) [17] introduced the concept of transfer learning. BERT was pre-trained on a large corpus of generic, unlabeled data (e.g., Wikipedia), and then fine-tuned on smaller, labeled downstream datasets to generalize across various tasks. One of the pre-training objectives used in BERT [17] is masked language modeling (MLM). In this objective, a percentage of the tokens of each sequence is masked randomly, and the model is optimized to correctly predict these masked tokens.
Although the transfer learning scheme employed in BERT has yielded promising results for numerous tasks [17, 18], its effectiveness is limited when applied to downstream tasks that fall outside the domain of the pre-training corpus. For example, a model trained on Wikipedia will not be able to capture the nuances of medical or legal languages. To address this, specialized models have been pre-trained on domain-specific corpus rather than general-purpose [19, 20]. However, such specialized models require a large and diverse corpus from the specific domain to accurately capture its nuances. When a sufficient domain-specific corpus is not available, an intermediate step, known as domain adaptation (DA), is often performed. In this process, the generic model is further trained on the available unlabeled domain-specific corpus [20]. The further training step in DA is expected to update the model’s weights, integrating knowledge from the desired domain alongside its already established knowledge from the pre-training data.
Molecular transformer models used for property prediction tasks have predominantly followed the pre-training $\rightarrow$ fine-tuning scheme, with few exceptions. For instance, K-BERT [21] was initially pre-trained on a dataset that lacked chiral information. To address this limitation, the model was later further-trained on a chirality dataset with an additional chirality classification objective. Many models have employed generalizable domain-specific objectives during pre-training. For example, MolBERT [1] and ChemBERTa-2 [22] trained their models to predict around 210 physicochemical properties for each molecule in the pre-training dataset (referred to as PhysChemPred and MTR in the two manuscripts, respectively). K-BERT [21] also trained their model to predict features per atom and a structural vector of the molecule calculated using the MACCS structural keys algorithm. Another widely used domain-specific objective is contrastive learning (CL) [23], which has been adopted by models such as Chemformer [24], MolBERT [1], and Transformer-CNN [25] using SMILES sequence augmentation. While domain-relevant objectives are argued to improve performance [1, 22], interpretability [21], and model stability [26], they can be computationally expensive to implement on large pre-training datasets [22]. To date, models incorporating domain-specific objectives during the computationally demanding pre-training step.
Although current molecular transformer models have been pre-trained on millions to billions of molecules, investigations have shown that increasing the pre-training dataset size does not consistently lead to improved predictions of molecular properties [15]. Several studies have explored the impact of increasing pre-training dataset size [2, 22, 27, 28], but this has not yet been done exhaustively, and these studies often lack distribution or significance analysis, which is crucial for ensuring the robustness. For example, pre-training on molecules combined from different databases, like ZINC and PubChem as in MolFormer [2] or ZINC, ChEMBL, and PubChem as in Chen et al. [28], did not show noticeable differences in performance compared to using a single database. Additionally, it has recently been demonstrated in the field of material science that large databases often contain a substantial amount of redundant information [29], while for the protein language modeling field, a small but diverse pre-training dataset was sufficient to improve performance on downstream analysis [30].
These observations [15, 29, 30] let us hypothesize that a current limitation in witnessing the power of transformer models in the molecular property area might be the redundancy in the pre-training dataset and its limitedness in capturing nuances and patterns that are causative for the measured target value of the downstream molecules. In this work, we therefore try to answer the following research questions:
  1. 1.
    How does increasing pre-training dataset size affect molecular property prediction?
  2. 2.
    How does domain adaptation with different objectives affect molecular property prediction?
  3. 3.
    What is the most efficient training approach without compromising performance?
  4. 4.
    How does domain adaptation compare to currently existing transformer and baseline models?
We hypothesize that performing domain adaptation (DA), as represented by the further-training step, will introduce the model to more relevant datapoints, thereby improving its performance. To perform DA, the domain-specific molecules are selected as the new unlabeled data for further training. Additionally, we propose that using domain-specific objectives (e.g., learning physicochemical properties) during this step will enhance performance due to increased chemical awareness. Employing domain-specific objectives in the DA step can also provide computational efficiency since the size of the DA data is significantly smaller than the pre-training data. To this end, we experimented with two objectives: multi-task regression (MTR) of physicochemical properties, and contrastive learning (CL) of different SMILES representations of the same molecules. In this scheme, the model is pre-trained on a large database of molecules using the MLM objective and then further-trained on the molecules from the downstream tasks using either MLM, MTR, or CL.
Our findings from the six datasets of the ADME benchmark [5] indicate that increasing pre-training dataset size was either not beneficial or only helpful up to a certain fraction of the dataset (i.e., $\ll 1$M molecules). We also observed that applying domain adaptation using the MTR objective significantly improved the model’s performance for solubility, permeability, and microsomal stability endpoints by showing P-values < 0.01, as well as noticeably increasing the average $R^{2}$ by more than 0.1 unit (i.e., large effect size). Additionally, transformer models can be trained much more efficiently than existing models while preserving or even enhancing predictive power. Finally, all examined transformer models performed similarly to a basic RF model with default settings on a pre-defined physicochemical properties as a molecular representation on the examined datasets.
We believe that molecular transformer models are yet to be optimized to address the specific needs and nuances of molecular data, such as ensuring diversity in the pre-training dataset, selecting the most suitable training objectives, and understanding the relation between upstream and downstream datasets. Therefore, we expect these models to surpass current baselines when such measures are taken into account.
2 Datasets and Preprocessing
In this work, we pre-train a transformer model on a dataset sampled from a large-scale general-purpose library of molecules, namely GuacaMol [18]. We then evaluate our models on a benchmark dataset that explores ADME properties (Absorption, distribution, metabolism, excretion) [5]. In our approach, datasets are utilized in three stages of model training: pre-training, domain adaptation, and evaluation. The GuacaMol dataset is exclusively used for pre-training, the unlabeled ADME dataset for domain adaptation, and the labeled ADME dataset version for evaluation. In this section, we provide a brief description of each dataset.
2.1 Pre-training Dataset: GuacaMol
We pre-train the models on the GuacaMol dataset [18]. This dataset is a subset of $\sim$1.3M molecules sampled from the ChEMBL database [31], which contains molecules that have been synthesized and assayed with respect to different biological endpoints. The GuacaMol subset was filtered to benchmark generative models in multiple aspects such as molecular generation validity, or similarity. To this end, the ChEMBL database was filtered so that the selected molecules were different from a holdout set of 10 already marketed drugs. The filtering process was done by removing all molecules that have a similarity above $0.323$ to these 10 drugs measured as Tanimoto similarity using ECFP4 fingerprints. The molecules were also filtered by size, i.e., to contain between 5 and 100 atoms, and by element, i.e., to be part of this atom list: H, B, C, N, O, F, Si, P, S, Cl, Se, Br, and I. Furthermore, they were standardized, i.e., salts were removed and charges neutralized.
2.1.1 Pre-training dataset selection
Besides training a model on the full training split of the GuacaMol [18] dataset (i.e., $\sim 1.3$M molecules), we investigate three more models trained on 0% (i.e., randomly initialized model with no pre-training), 30%, and 60% of the datasets. To select the 30% and 60% subsets, the data was first clustered using BitBirch [32] based on Morgan fingerprints (radius 2, 2084 bits) and tanimoto similarity, then selections were made to ensure the same cluster distribution per subset. RDKit [33] was used to calculate fingerprints and similarity, and the selection procedure was adopted from talktorial T005 from TeachOpenCADD [34].
2.2 Domain Adaptation and Downstream Datasets: ADME benchmark
The domain adaptation (DA) step is introduced to further train the pre-trained model on a smaller domain-specific dataset. In our analysis, the molecules of the downstream dataset are passed to the pre-trained model without labels for DA. Later, in the MPP task the ADME dataset molecules including their labels are used for evaluation. The full benchmark dataset contains $\sim 3.5$K molecules, however, each of the four ADME endpoints was tested on a subset only. Table 1 shows the number of molecules tested per endpoint. In the following, we will detail the different endpoints.
# Molecules # Molecules without censored labels
Solubility 2173 2173
Permeability 2642 2642
HLM 3087 2129
RLM 3054 2708
hPPB 194 194
rPPB 168 168
Table 1: Number of molecules per endpoint with and without censored labels. The numbers in bold refer to the datasets that changed in size after removing censored labels.
Solubility is a vital property for understanding the uptake and distribution of a molecule within living organisms, which determines a molecule’s efficacy and usability [35]. Solubility assays inform about maximum amount of a molecule to dissolve in a solution, either water or a physiologically relevant medium, under certain pH and temperature [36]. Fang et. al. [5] measured the solubility of $\sim 2.2$K molecules in phosphate buffered saline (PBS) at pH 6.8 following the protocol in Kestranek et. al. [37]. The reported measuring unit is $\mu g/mL$. Figure1 shows the distribution of the measurements. The data has a bimodal distribution with a major peak at $\sim 1.5$ log($\mu g/mL$) ($\sim 31.62$ $\mu g/mL$) and another small peak at $\sim-0.5$ log($\mu g/mL$) ($\sim 0.316$ $\mu g/mL$).
Permeability assays determine the rate at which a molecule passes through a biological membrane by measuring the ratio between movement to and from the bloodstream [38, 39]. Fang et. al. uses the MDR1 transfected MDCK cell lines to measure this efflux ratio [5]. MDR1 (also known as P-glycoprotein or P-gp receptor) is present in various tissues, such as intestines and blood-brain barrier, making it a good candidate to assess permeability [40, 38, 39]. Figure 1 shows the distribution of the $\sim 2.6$K molecule’s efflux ratio, where $\sim 50\%$ of the data is concentrated at $\sim-0.2$ Log unit (an efflux ratio of $\sim 0.6$).
Microsomal stability is used to measure the clearance of a molecule from the body, specifically, intrinsic clearance $CL_{int}$ [41]. Intrinsic clearance quantifies the volume of incubation medium from which a molecule is cleared per time unit per microsomal weight unit. In Fang et. al. [5], the measuring unit is mL/min/kg presented for two datasets: human liver microsomes (HLM) and rat liver microsomes (RLM). Figure 1 shows the data summary of the two datasets with HLM spanning around three log units of mL/min/kg (0.6 - 3.4), while RLM spans around four log units (1 - 4). The HLM dataset has a sharp peak at $\sim 0.67$ ($\sim 4.7$ mL/min/kg). Interestingly, the boxplot and the scatter plot of the individual data points show that $\sim 27\%$ of the data has the same target value. This observation corresponds to what is called "censored labels", where experimental values can be recorded only up to a certain threshold, and every observation beyond this threshold is set to a constant value [42]. When training a regression model to predict continuous values, this censoring is introducing noise to the data. The same censoring problem is seen for the RLM dataset, although less severely as only $\sim 9$% of the data is censored. While the censored regression problem has been investigated in fields like survival analysis, its resolution in MPP is still in progress [42]. In this work, we handle the censored label by removing the molecules with censored labels from the evaluation step to reduce the noise. However, we keep the molecules during the DA step as they still provide information for further training. After removing these values, the number of labeled molecules becomes $\sim 2.1$K and $\sim 2.7$K for HLM and RLM datasets, respectively. Suppl. Figure S1 and Table 1 show the new distribution and summary of the two datasets. The HLM dataset now spans the range 0.7 to 3 log units while RLM spans the same range (1 to 4 log units).
Plasma protein binding assays measure the amount of a compound that has not been captured by the plasma proteins in the blood. This assays is important because it is believed that the percentage of free molecules in the blood is responsible for a compound’s activity and efficacy. Therefore, knowing the percentage of free molecules can also guide other assays like permeability and clarence [43]. Fang et. al reported the percentage of unbound molecules in the blood in both human (hPPB) and rats (rPPB). Both datasets contained very few molecules ($<200$ molecules) and the measurements span the range -2 to 2 Log unit (0.01 - 100% unbound molecules), see Figure1.
Figure 1: Data summary and distributions of the six investigated endpoint.
2.3 Dataset Splitting
For pre-training, we used the specified split of the GuacaMol dataset into train and test splits with an 80:20 ratio. For domain adaptation, all endpoint molecules were used for further-training. Finally, 5-fold cross validation (CV) was performed for evaluating the predictions on the labeled molecules. We report two types of splitting for CV: random splitting, and cluster-based splitting using DataSAIL [44] with default settings. To avoid data leakage, DataSAIL splits data by minimizing the similarity between any two splits formulated as a Binary Linear Program (BLP) to optimize the solution. 5-fold cluster-based splitting was possible for all datasets except for hPPB, which could only be split 3-fold.
3 Methods
In this work, we trained a BERT-like model with the simple MLM objective using the GuacaMol [18] dataset followed by domain adaptation on the ADME benchmark [5] using either MLM or a domain-specific objective, i.e., MTR and CL (Figure 2). We trained multiple models using different sizes of pre-training datasets to further assess the relationship between pre-training dataset size and downstream performance. In the following, we will detail the architecture and training of our models. We describe each training objective, explain the method for increasing the pre-training dataset size, list the literature models used for benchmarking, and finally, outline the evaluation process. Code, data, and analysis can be found at our github repo https://github.com/uds-lsv/domain-adaptation-molecular-transformers
Figure 2: An overview of this research’s workflow. Transformer models are trained by pre-training on generic large unlabeled datasets using one or more objectives (step 1), followed by fine-tuning on labeled datasets (step 2). Domain adaptation is an optional intermediate step that resembles pre-training, but can be done on much smaller unlabeled dataset (step 1.5).
3.1 Model Architecture and Training
In this work, we use the SMILES representations of the molecules as input to our models. We used the SMILES strings as provided by both publications without further standardization [18, 5]. For the transformer model, we use the same architecture as the original BERT [17] model with 12 layers, 12 attention heads, and 768-dimensional embeddings, which yields $\sim 89M$ parameters. We use a maximum sequence length of 128 tokens and a vocabulary size of 4096 tokens. We also used absolute positional embeddings and the same tokenizer (i.e., WordPiece [45, 46]) as BERT. We trained in mixed precision using bf16.
For pre-training, we use the MLM objective (explained below) with a batch size of 16 for 20 epochs (similar to MolBERT [1]). We used Adam [47] with a learning rate of $3e^{-5}$ and a linear scheduler with 10% warm-up. The same hyperparameters were used for domain adaptation, however, the objectives varied between MLM, MTR, and CL (explained below).
For evaluation, we used either the pre-trained model or the domain-adapted model to generate fixed embeddings using the CLS token (i.e., the first token fixed in front of any sequence) for each molecule in the test set. These embeddings were used as representations of the molecules and were used as an input for a random forest (RF) regressor model from sklearn [48] (default settings were used).
3.2 Training objectives
The transformer models learn by optimizing a self-supervised training objective like masked language modeling (MLM). In this work, we used MLM as an objective for pre-training, and explored it for domain adaptation as well. Multi-task regression (MTR) was used for domain adaptation and as a pre-training objective for one of the experiments. Contrastive learning (CL) was used for domain adaptation only. below, we briefly explain each of these objectives. For all objectives, we used a batch size of 16 with a learning rate of $3e^{-5}$ for 20 epochs. The maximum input length was 128 and mean pooling of the tokens was applied to extract a molecule’s representation.
Masked Language Modeling (MLM)
In the MLM objective, 15% of the input SMILES tokens are randomly masked in each molecule. The language model then outputs the probability distribution over all possible tokens in the vocabulary for each masked token. MLM is quantified by minimizing the cross-entropy loss:
$L_{MLM}=-\sum_{c=1}^{M}y_{o,c}\log(p_{o,c})$ (1)
Where $M$ corresponds to the total number of classes (i.e., the vocab size), $y_{o,c}$ is the true class of token o, and $p_{o,c}$ is the predicted class.
Multi-Task Regression (MTR)
The MTR objective is independent of the linguistic structure of the SMILES. This objective involves the simultaneous prediction of a vector of $210$ real-valued physicochemical properties of the input molecule. In our work, we use the RDKit [33] framework to calculate the physicochemical descriptors of the pre-training dataset molecules, we normalize the values using the standard scaling, and predict properties using 2 layer MLP with relu activation functions and dropout of 0.1 on the first token. The model uses multi-task mean squared error loss
$L_{MTR}=\sum_{i=1}^{N}\sum_{j=1}^{D}(p_{ij}-y_{ij})^{2}$ (2)
where $D$ is the $210$-dimensional chemical descriptors and $N$ is the number of training samples, $p_{ij}$ is the predicted value and $y_{ij}$ is the true value. For smooth convergence, we use mean and standard deviation to normalize each descriptor as a pre-processing step.
Contrastive Learning (CL)
The CL objective utilizes the nuances of the SMILES sequence since multiple sequences can represent the same molecules, a process called enumeration. A canonical SMILES sequence per molecule can be generated by fixing the starting and branching procedure of linearizing the molecular graph. We employ CL by using the multiple negatives ranking loss [49] which takes SMILES triples. Each triple consists of a canonical SMILES, an enumerated sequence of the same molecule, and the SMILES of a random molecule from the dataset as negative example. The encoder is based on a single BERT model, in which the latent SMILES representations of the canonical and enumerated SMILES are pulled together, whereas the latent representation of the negative SMILES are pushed away from the latent representations of the other two SMILES simultaneously. The contrastive loss is expressed as follows:
$L(r_{c},r_{e},r_{n})=\frac{1}{K}\sum_{i=1}^{K}\left[{\rm sim}(r_{c}^{(i)},r_{e}^{(i)})-\log\sum_{j=1}^{K}e^{{\rm sim}(r_{c}^{(i)},r_{n}^{(j)})}\right]$ (3)
[…]
3.3 Models benchmarking
To identify the strengths and weaknesses of our model, we compare the performance in downstream task prediction when training a RF model using embeddings of our models versus embeddings of two other transformer models, MolBERT [1] and MolFormer [2], as well as the physicochemical properties calculated by RDKit [33]. We used the provided MolBERT [1] implementation and the MolFormer’s huggingface [50] implementation to extract the embeddings. The same RF model with default settings as provided by scikit-learn [48] was used to predict the molecular properties using the embeddings of our models, MolBERT, MolFormer, and the physicochemical properties. Table 2 shows the differences between all models examined. The RF + PhysChem model is considered as a simple and transparent baseline model to compare against the remaining models.
Model Pre-training Domain Adaptation # Parameters
# Molecules ($\sim$) Objectives # Molecules ($\sim$) Objectives
Ours 0, 390K, MLM or MTR 170 – 3K MLM, MTR, $\sim$89M
780K, 1.3M or CL
MolBERT 1.3M MLM, MTR, - - $\sim$85M
SMILES-Eq
MolFormer 100M MLM - - $\sim$85M
RF + Physchem - - - - -
Table 2: An overview of the benchmarked models explaining their training decision in terms of pre-training dataset size, objectives and number of trainable parameters. MLM: masked language modeling, MTR: multi-task regression of physicochemical properties, SMILES-Eq: predicting whether two SMILES strings correspond to the same molecule, and CL: contrastive learning of the different representations of the SMILES string.
3.4 Evaluation metrices
During training, we normalize the response values using the standard scaling, for the evaluation we transform predictions back to the original values. We report performances using the coefficient of determination ($R^{2}$, equation 6), which measures the goodness of fit by comparing the model’s residual sum of squares (RSS, equation 4) (i.e., true value $y_{i}$ to predicted value $\hat{y_{i}}$) to the total sum of squares (TSS, equation 5) (i.e., true value $y_{i}$ to the mean value $\bar{y}$). A value of zero means that the model is performing as good as a model that only predicts the mean value of a dataset. A value $>0$ corresponds to a model that gives better prediction than such mean-only model, while a value $<0$ corresponds to a model worse than the mean-only model. The maximum value of the metric is 1 (i.e., the model predicted each value correctly) and the lowest value is $\infty$- ∞. The sklearn package [48] was used to calculate the $R^{2}$ metric.
$\text{RSS}=\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (4)
$\text{TSS}=\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}$ (5)
$R^{2}=1-\frac{\text{RSS}}{\text{TSS}}$ (6)
In our analysis, we show the $R^{2}$ of the five folds as either a boxplot or an error barplot using the standard error of means (SEM, equation 7), which assesses the variability of model estimates using the sample standard error $\sigma$ and the sample size $n$.
$\text{SEM}=\frac{\sigma}{\sqrt{n}}$ (7)
For significance analysis, we report P-values comparing the different models using the non-parametric Mann-Whitney U test (also known as Wilcoxon rank-sum test) [51]. This test does not require the data to be normally distributed, and is calculated by ranking the values in two distributions to determine if there is a significant difference in the sum of ranking of each distribution. The P-value of multiple comparisons is corrected using the Holm’s method [52] to account for the problem of increased probability of false positives [53]. The scikit-posthoc package [54] was used to calculate the corrected P-values, and a P-value $<0.05$ was considered significant.
4 Results and discussion
In this section, we discuss our findings on the effect of increasing pre-training dataset size, the benefit of using domain adaptation (DA), and how the most efficient and performant setup compares to models from the literature.
4.1 Pre-training improves performance on downstream endpoints up to a limit
Frist, we compare the model performances with varying pre-training dataset sizes. Figure 3 shows the performance of a model pre-trained with different increasing percentages of the GuacaMol [18] dataset molecules from 0% (i.e., randomly initialized wights) to 100%. The boxplots for the different pre-training subsets show that increasing size improves performance until 30% of the data as can be seen for solubility, permeability, HLM, and RLM. Significance analysis between the 0% and 30% setups supports these observations for solubility, permeability, and RLM (P-value $<0.05$, Suppl. Figure S2). However, models trained on further increments in pre-training data show similar and overlapping boxplots to the 30% setup (P-value >0.05)>0.05)> 0.05 ), Suppl. FigureS3) .
Figure 3: The effect of increasing pre-training dataset size using DataSAIL splitting. 0% corresponds to a randomly initialized model with no pre-training and 100% correspond to the $\sim 1.3$M molecules of the GuacaMol dataset.
The hPPB and rPPB datasets pose a challenge to draw insights from due to the very low size of the datasets (i.e, $<200$ molecules), therefore, concluding a positive or negative result is not well supported.
Examining the values of the $R^{2}$ metric, one can observe that the permeability dataset showed high $R^{2}$ values as most of the values lie above 0.2, while most of the $R^{2}$ values for the rest of the datasets were below 0.15 (Figure 3). As a reminder, an $R^{2}$ value of 0 corresponds to a model as predictive as a model that predicts only the mean value, the upper bound is 1, which indicates a model significantly better than predicting the mean, and the lower bound is $\infty$- ∞, corresponding to a model worse than predicting the mean. Therefore, having most of the $R^{2}$ values at such low range imply that the different models still as naive as a mean-predictor despite the significance difference between the different setups.
Suppl. Figures S5 and S6 show comparable results under the more relaxed random splitting scheme. In these figures, the same trend holds as the performance improves significantly when pre-training on 30% of the pre-training data compared to a randomly initialized model (P-value $<0.05$ for all endpoints except plasma protein binding, Suppl. Figure S4), however, no significant improvement is observed beyond 30%. Finally, even though the random split is known to overestimate a model’s performance due to possible data leakage [55, 4], the $R^{2}$ values are still lying below 0.2 for solubility, HLM, and RLM, which indicates that the models are still as naive as a mean-predictor even for this relaxed split.
The lack of diversity can explain the limitedness of pre-training.
To better understand these findings, we explored the composition of the individual datasets in terms of molecular scaffold analysis. Table 3 shows the percentage of unique Murcko scaffolds in both pre-training (GuacaMol) and each endpoint datasets. The pre-training dataset consists of $\sim 1.3$M molecules and $\sim 407$K unique scaffolds. The percentage of unique scaffolds compared to the size of the dataset does not exceed 32%, which suggests that the pre-training dataset is low in diversity. This might explain the reason for the limited performance gain when using the full dataset. This is also inline with the findings from the neighboring field of material science where most databases were found to be redundant [29]. As a reminder, we select the different pre-training subsets by sampling from the different the dataset after clustering it based on tanimoto similarity and Morgan fingerprints using the BitBirch algorithm [32]. Therefore, most of the diversity of the dataset has been selected in the first 30% subset. Additionally, the table shows that the percentage of unique scaffolds per endpoint is extremely high (i.e., from 87% to 91%), while the percentage of common scaffolds with the pre-training dataset were between 31% and 36%. This suggests that the downstream molecules are too diverse on their own, and quite diverse from the pre-training dataset.
As a reminder, the objective used for pre-training is the masked language modeling (MLM), which exclusively learns patterns from the molecular structure as expressed by the tokenization of the SMILES sequence. Therefore, the low structural diversity of the pre-training dataset means that the model will learn frequent patterns quickly but it will struggle to generalize to structurally different sequences -as observed in our case-. This suggests that the MLM objective is not the most appropriate objective given the state of the datasets, which can also explain the low predictive power of the models.
% of Unique Murcko scaffolds ($\sim$) % of common scaffolds ($\sim$)
(GuacaMol $\cap$ endpoint)
GuacaMol 32
Solubility 87 34
Permeability 88 31
HLM 90 32
RLM 88 31
hPPB 91 32
rPPB 90 36
Table 3: Percentage unique Murcko scaffolds in pre-training and downstream datasets, as well as percentage of the common scaffolds between each endpoint moleculs and the pre-training dataset.
Conclusion:
Pre-training improved downstream endpoint predictions using 30% of the data. However, the usage of the full dataset did not provide significant improvements. This can be attributed to the low diversity of molecules — in form of Murcko scaffolds — in the pre-training dataset. Thus, adding more data does not necessarily introduce new patterns for the model to recognize. Generally, the model performances quantified by the $R^{2}$ metric were on the lower range for both the strict DataSAIL splitting and the relaxed random splitting, suggesting that the models are not much better than a simple model that only predicts the mean value. Further improvements in terms of pre-training dataset diversity and objective selection can help improve the models’ performance.
4.2 Domain adaptation improves performance significantly
Here, we compare domain adaptation with different objectives to the models without pre-training across the different pre-training dataset sizes. Figure 4 shows noticeable improvement with DA using the MTR objective for Solubility, Permeability, HLM, and RLM for all different pre-training setups. The CL objective also shows noticeable improvement compared to the No DA model, however, less prominent across the different pre-training setups compared to the MTR objective. Finally, DA using the MLM objective was not helpful as it performed comparable to the models without domain adaptation in all datasets. Significance analysis further corroborate these observations (Figure 5) as the DA models using MTR is showing P-values $<0.05$ for all pre-training setups for solubility, permeability, HLM and RLM, while DA using the CL objective shows P-values $<0.05$ for at least two of the different pre-training setups for the same datasets. DA using the MLM objective shows no significant difference to the No DA models for all datasets (P-value > 0.2).
Suppl. Figures S7 and S8 show comparable results for the random split experiments as DA using MTR and CL are both significantly better than the models with no domain adaptation for all datasets.
Figure 4: The effect of domain adaptation with different sizes of pre-training dataset using DataSAIL splitting. SE corresponds to the standard error of means. 0% corresponds to a randomly initialized model with no pre-training, and 100% corresponds to pre-training on the $\sim 1.3$M molecules of the GuacaMol dataset. No DA: a model with no domain adaptation, MTR: domain adaptation using multi-task regression of physicochemical properties, CL: domain adaptation using contrastive learning of the different representations of the SMILES string, and MLM: domain adaptation using masked language modeling.
Figure 5: Significance analysis of domain adaptation with different sizes of pre-training dataset using DataSAIL splitting as shown in Figure 4
Basic physicochemical properties might explain why MTR works.
DA analysis showed that MLM is not improving the model’s performance, CL is improving it to some extent, and MTR is improving it greatly. Here, we explain this behavior by touching back on the diversity feature of both datasets as well as the distribution of some basic physicochemical properties of the endpoint molecules. Table 3 shows that the molecules in the downstream enpoints are extremely diverse in terms unique Murcko scaffolds. Since MLM is an objectives that learns the sequence only, it will struggle to capture consistent patterns of the endpoint molecules, which can explain the weak performance of this objective during DA. The CL objective is also used as a structure-based objective like MLM, therefore, it is affected by the limitation in the structural composition of the dataset. However, the model sees multiple SMILES representations of the same molecule, which helps it capture patterns more consistently than MLM. This can explain the improved performance of the CL objective during DA compared to no DA at all or DA using the MLM objective. Suppl. Figures S12, S13, S14, S15, S16 and S17 illustrate the distribution of the molecules tested per endpoint for six physicochemical properties: molecular weight (MW), topological polar surface area (TPSA), the partition coefficient (LogP), number of hydrogen bond donors (HBD), hydrogen bond acceptors (HBA), and rotatable bonds. While the majority of molecules per endpoint possess unique Murcko scaffolds, they tend to cluster in a narrow range of histogram bins in the physicochemical property analysis (i.e., they share similar physicochemical profile). This explains why MTR is the best performing DA objective (Figures 4, S7, 5, and S8) as the molecular space is already easier to cluster by the physicochemical properties.
Conclusion:
Domain adaptation using the MTR objective significantly improved performance for three out of the four endpoints (solubility, permeability, and microsomal stability) compared to the models without domain adaptation. DA using the MLM objective did not yield significant improvements to the pre-trained-only models while the CL objective showed significant improvement in some setups. The MLM objective’s limited performance may be attributed to the high diversity in Murcko scaffolds for the downstream dataset. In contrast, the MTR objective was effective because the downstream data shared similar physicochemical profiles despite structural differences. Finally, the CL objective provided intermediate performance, improving upon the MLM objective by enabling the model to learn more consistent molecular representations.
4.3 MTR is better for pre-training, but best for domain adaptation.
Given the MTR objective’s promising performance during domain adaptation, we further explored if MTR would also outperform MLM during pre-training (Figure 6). The MTR objective consistently outperformed the MLM objective across all pre-training setups for solubility, permeability, HLM, and RLM. The same trend was observed with random splitting (Suppl. Figure S18). Additionally, the same trend of no significant improvement beyond 30% of the pre-training dataset was again observed when MTR was the pre-training objective (Suppl. Figures S19, S20, S21 and S22).
Figure 6: Performance comparison between the models pre-trained with MLM (blue) vs MTR (orange)objectives. DataSAIL splitting was used. 0% corresponds to a randomly initialized model with no pre-training, and 100% corresponds to pre-training on the $\sim 1.3$M molecules of the GuacaMol dataset.
Figure 7: comparing pre-training and domain adaptation objectives to identify the best approach. the models names follow the scheme of pre-training objective_pre-training dataset size_domain adaptation objective. DataSAIL splitting was performed.
Given that MTR is generally more expensive to optimize than MLM [22], and pre-training is inherently more computationally expensive than domain adaptation, it is crucial to allocate the best objective at the right stage to achieve optimal predictive power efficiently. Since pre-training with only 30% of the data consistently provides the best downstream performance, we focused on models with this setup. In the following, the performance of the models pre-trained with either MLM or MTR objective, and with either no domain adaptation or with MTR as a domain adaptation objective is investigated (Figure 7 and Suppl. Figure S23 for DataSAIL and random splitting, respectively). The figures show that models pre-trained with MLM and domain adapted with MTR (MLM_30%_MTR) perform similarly to those pre-trained with MTR, with or without domain adaptation. This is further supported by significance analysis (P-values $>$ 0.05, Suppl. Figures S25 and S24 for DataSAIL and random splitting, respectively). Given that MLM_30%_MTR is the most efficient setup without compromising performance, we conclude that MTR is best for domain adaptation rather than pre-training for the currently analyzed endpoints.
Basic physicochemical properties distributions of the upstream and downstream datasets can help understand this observation.
Our analysis demonstrated that the MTR objective for pre-training outperforms the MLM objective for predicting downstream properties. However, it did not improve performance when it was used in both pre-training and domain adaptation compared to it being used for domain adaptation only. To understand this behavior, we examined the distribution of basic physicochemical properties between the two datasets which showed that the downstream dataset spans similar distributions to the upstream dataset (Figure 8). Given that the two datasets share minimal molecules and exhibit significant diversity in Murcko scaffolds (Table 3), the MLM objective then proves insufficient for this scenario and the MTR objective is proven more helpful for pre-training. In both pre-training and domain adaptation, the upstream dataset spans much wider ranges than the downstream dataset (Figure 8), indicating that, although similar, the distributions still show many out-of-distribution molecules and physicochemical properties. Although the significance analysis indicates that none of the three models (MLM_30%_MTR, MTR_30%No DA, and MTR_30%_MTR) performs significantly better than the others, a visual inspection of the boxplots reveals that the distribution for the MLM_30%_MTR model is more compact, and consistently showing higher mean and median values. Thia observations leads us to suspect that the domain adaptation step for the model pre-trained with MTR is constrained by the patterns learned during pre-training. In contrast, domain adaption for models pre-trained with MLM appears to facilitate easier learning from the downstream dataset. This can help explain why MTR is more suited for domain adaptation rather than pre-training.
Figure 8: The distribution of basic physicochemical properties between the pre-training dataset (GuacaMol) and all the molecules in the downstream benchmark. The six physicochemical properties are molecular weight (MW), topological polar surface area (TPSA), the partition coefficient (LogP), number of hydrogen bond donors (HBD), hydrogen bond acceptors (HBA), and rotatable bonds.
Conclusion:
The results confirm that the MTR objective is more suitable than the MLM objective for the tested pre-training and downstream datasets. However, the added benefit of pre-training also stops after 30% of the dataset. Finally, although MTR is better as a pre-training objective, it is more efficient (without compromising performance) to use it during domain adaptation only and to employ the MLM objective during pre-training.
4.4 Our efficient transformer model rivals more complex models.
As examined so far, domain adaptation with the MTR objective using fewer samples ($\leq$ 3K molecules) is competitive against extensive pre-training beyond 400K molecules. The optimal model (i.e., MLM_30%_MTR) combines MLM for pre-training and MTR for domain adaptation, matching the predictive power of MTR for both stages. We compare the performance of the MLM_30%_MTR model to two other transformer models, MolBERT [1] and MolFormer [2]. MolBERT was pre-trained on $\sim$1.3M molecules using the GuacaMol dataset and three objectives were performed simultaneously during pre-training, MLM, MTR, and predicting whether two SMILES sequences correspond to the same molecule (SMILES-Eq). MolFormer was trained on 100M molecules compiled from the ZINC and PubChem databases while using only the MLM objective. A baseline is also shown for an RF model with default settings as provided by the scikit-learn library and trained on the physicochemical properties of the molecules.
The results in Figure 9 indicate that our model performs similar to MolBERT in all datasets, as evidenced by the close values of $R^{2}$ in the boxplots, and outperforms MolFormer for all datasets except rPPB. However, the significance analysis reveals that the three models perform similarly for all datasets (P-values $>0.05$), except for permeability, where both our model and MolBERT significantly outperform MolFormer (P-value = 0.048, Figure 10).
While our efficient model demonstrates competitive performance compared to more complex transformer models, both our model and MolBERT perform similarly to the simple RF + PhysChem model across all datasets. MolFormer also achieves similar results to the RF + PhysChem model for all datasets except permeability and RLM, where the simple RF model performs significantly better (P-value = 0.048).
Similar trends hold for all compared models when the random splitting scheme is used (Suppl. Figures S10 and S11)
Figure 9: Performance of the most efficient model from our analysis to models from the literature. MLM_30%_MTR corresponds to a transformer model pre-trained with 30% of the GuacaMol dataset using the MLM objective then domain adapted on the corresponding endpoint using the MTR objective. MolBERT [1] and MolFormer [2] are two transformer models from the literature. RF + PhysChem corresponds to a model trained on the physicochemical properties calculated by the RDKit package. The results are shown for the DataSAIL splitting scheme.
Figure 10: Significance analysis of the benchmarking shown in figure 9.
Yet opportunities for improvement remain
Our model that was pre-trained on fewer than $400$K molecules using the MLM objective and domain adapted on fewer than $3$K molecules with the MTR objective rivals the more complex and expensive MolBERT[1] and MolFormer[2] models. This suggests that current transformer models may be over-engineered and not optimized for the underlying chemical space. We already showed that the GuacaMol dataset is redundant, meaning it contains many very similar molecules, and is quite dissimilar to the endpoints molecules in terms of Murcko scaffolds, which could be limiting the benefit of pre-training. This affects both our models and MolBERT as both were pre-trained on the same dataset. MolFormer has been trained on 100M molecules combined from both ZINC and PubChem, therefore, spanning bigger and different chemical space than GuacaMol. All three transformer models have similar numbers of trainable parameters (i.e., $\sim 90$M), but are pre-trained on different magnitudes of molecules ($\sim 400$K for our model, $\sim 1.3$M for MolBERT, and $\sim 100$M for MolFormer). Scaling analysis of neural networks already suggests that there should be a fixed relationship between increasing data set size and the number of parameters, i.e. scaling laws, to achieve the desired improvement from such increases [56, 57]. Furthermore, a scaling experiment for molecular transformer models showed that molecular data exhibit a resolution-limited regime [58]. This regime suggests that a neural network trained on such data will need either a well-refined data set or a deeper neural network to capture enough nuances of the data [59]. Further analysis can investigate the best practices of model architecture and proper coverage of the chemical space.
Finally, a basic random forest (RF) model trained on physicochemical properties also showed competitive results to the three transformer models. We already showed that physicochemical properties are better descriptors of current downstream molecules than the structure alone. This is also highlighted when the same RF model was trained on the structural Morgan fingerprints, and it showed significantly worse performance than the RF model with physicochemical properties (Suppl. Figures S26, S27, S28 and S29). Our model, MolBERT and the RF models trained on the pyhsicochemical properties are outperforming models relying on structure-based representations like using Morgan fingerprints as input for RF or relying solely on MLM as in MolFormer. This shows that the main force behind the models’ performance is the molecular descriptors rather than the model architecture. Therefore, the current implementations of the transformer models is still limited in capturing fine-grained information about the molecules. Further investigations like probing and explainability can help identify the weak areas of the model’s learning, therefore, finding ways to further improve it.
Conclusion:
The results show that training a competitive transformer model can be done as efficiently as pre-training on $<400$K molecules with a simple objective like MLM, and domain adaptation on a couple of thousands of endpoint molecules using the MTR objective. Such a model can perform similarly to a highly domain-oriented model like MolBERT (i.e., using multiple domain objectives during pre-training) and better than a model pre-trained with a huge generic corpus like MolFormer (i.e., trained on 100M molecules from both ZINC and PubChem). However, the results also show that these three models, regardless of their complexity, are still on par with a basic RF model with default settings using physicochemical properties as molecular representation. We showed multiple reasons that can explain these observations, and we believe that the transformer models still have a large room for improvement to better represent molecular data and predict properties.
5 Conclusion
In this study, we trained a transformer model using either the MLM or the MTR objectives to predict six ADME endpoints (solubility, permeability, microsomal stability in both humans and rats, and plasma protein binding in both humans and rats). We experimented with different setups of pre-training dataset sizes, starting from a randomly initialized model to a model trained on the full GucaMol dataset of 1.3M molecules. We proposed domain adaptation as an extra training step where the pre-trained model is further trained on the molecules from the downstream task itself without providing labels. We also investigated the MTR and CL objectives to understand how chemically-aware objectives perform compared to a lexical objective like MLM. We showed that training a transformer model can be performed very efficiently when taking into account the data diversity and the appropriate training objectives. The most efficient, yet performant, model setup was pre-training with $\sim 400$K molecules using the MLM objective and domain adaptation using the endpoint molecules ($<3$K) using the MTR objective. Our model competes against more complicated models like MolBERT [1] (pre-trained on 1.3M molecules using MLM, MTR, and SMILES-Eq) and MolFormer [2] (pre-trained on 100M molecules using the MLM objective).
In conclusion, we highlighted the potential for developing efficient and high-performing molecular transformer models. We also highlighted areas of improvement like investigating pre-training data quality or diversity, identifying which objectives better suited for capturing the nuances of a property endpoint, and interrogating the architectural choices of the model.
6 Data and code availability
All data, analysis and implementation code can be found in our github repo at https://github.com/uds-lsv/domain-adaptation-molecular-transformers
Supporting Information
Figure S1: Data summary and distributions of the six investigated endpoint after removing censored labels
Figure S2: Significance analysis of the 0% vs 30% setups only. DataSAIL splitting is used
Figure S3: Significance analysis of the increased pre-training dataset size using DataSAIL splitting
Figure S4: Significance analysis of the 0% vs 30% setups only. Random splitting is used
Figure S5: The effect of increasing pre-training dataset size using random splitting.
Figure S6: Significance analysis of the increased pre-training dataset size using random splitting
Figure S7: The effect of domain adaptation with different sizes of pre-training dataset using random splitting.
Figure S8: Significance analysis of domain adaptation with different sizes of pre-training dataset using random splitting.
Figure S9: Significance analysis of domain adaptation across different sizes of pre-training dataset under the MTR objective. random splitting was performed.
Figure S10: Performance of the most efficient model from our analysis to models from the literature. 30% + MTR corresponds to a transformer model pre-trained with 30% of the GuacaMol dataset then domain adapted on the corresponding endpoint using the MTR objective. MolBERT [1] and MolFormer [2] are two transformer models from the literature. RF + PhysChem corresponds to a model trained on the physicochemical properties calculated by the RDKit package. The results are shown for the random splitting scheme.
Figure S11: Significance analysis of the literature comparison.
Figure S12: Physicochemical properties distributions and data clusters of the Solubility endpoint
Figure S13: Physicochemical properties distributions and data clusters of the permeability endpoint
Figure S14: Physicochemical properties distributions and data clusters of the HLM endpoint
Figure S15: Physicochemical properties distributions and data clusters of the RLM endpoint
Figure S16: Physicochemical properties distributions and data clusters of the hPPB endpoint
Figure S17: Physicochemical properties distributions and data clusters of the rPPB endpoint
Figure S18: Performance comparison between the models pre-trained with MLM vs MTR objectives. Random splitting was used.
Figure S19: The effect of increasing pre-training dataset size for the MTR-pre-trained models using DataSAIL splitting.
Figure S20: The effect of increasing pre-training dataset size for the MTR-pre-trained models using Random splitting.
Figure S21: Significance analysis of the increased pre-training dataset size with the MTR objectives using DataSAIL splitting
Figure S22: Significance analysis of the increased pre-training dataset size with the MTR objectives using random splitting
Figure S23: comparing pre-training and domain adaptation objectives to identify the best approach. the models names follow the scheme of pre-training objective_pre-training dataset size_domain adaptation objective. For example, MLM_30%_No DA corresponds to a model pre-trained with the MLM objective on 30% of the pre-training dataset and without domain adaptation. Random splitting was used
Figure S24: Significance analysis of the different models from figure S23. Random splitting was used
Figure S25: Significance analysis of the different models from figure 7. DataSAIL splitting was used
Figure S26: Comparison of a random forest model trained with structural descriptors (i.e., Morgan fingerprints) and physicochemical descriptors. DataSAIL splitting is used.
Figure S27: Significance analysis of a random forest model trained with structural descriptors (i.e., Morgan fingerprints) and physicochemical descriptors (figure S26.
Figure S28: Comparison of a random forest model trained with structural descriptors (i.e., Morgan fingerprints) and physicochemical descriptors. random splitting is used.
Figure S29: Significance analysis of a random forest model trained with structural descriptors (i.e., Morgan fingerprints) and physicochemical descriptors (figure S28).